An electric heater is immersed in a well-stirred tank. Solvent at 19°C, which has a heat capacity of 2.1 kJ/(kg°C), is fed at a speed of 11 kg/h. The hot solvent is discharged at the same rate. The tank initially contains 180 kg of cold solvent at 10°C. The heat flow of the electric heater is 884 W.
Calculate the time for the temperature of the solvent being discharged to be 63°C (write the answer in hours and with two decimals)
Answer #1

put Tz= 63°6 63 = 156.785 - 146 € -4-657X1057 793-785 = +146 e-+-697x10-5t -1.697 x 10-5t e Hence 0.64 23 = Hence t = 26001.0